OpenStudy (anonymous):

is x^2 = 100 x={-10,10} or {10}

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

well (-10)^2 = 100 also, so both.

OpenStudy (saifoo.khan):

A.

OpenStudy (anonymous):

so how do you know if there are parentheses around the -10

OpenStudy (anonymous):

um...for just x^2 = 100. you assume both. the difference is are you squaring -10 or are you doing something like 0 - 10^2

OpenStudy (anonymous):

umm k

OpenStudy (slaaibak):

you know there's brackets, because x^2 = x * x Therefore -10 * -10 = (-10)^2

OpenStudy (anonymous):

"A negative times a negative equals a positive" 10 x 10 equals 100, and -10 x -10 also equals 100.

OpenStudy (slaaibak):

^\[A^2\]